OpenStudy (anonymous):

What is the value of f(5) for the function f(x) = 3x - 4?

OpenStudy (anonymous):

11

jimthompson5910 (jim_thompson5910):

f(x) = 3x - 4 f(5) = 3(5) - 4 ... replace every x with 5 f(5) = 15 - 4 f(5) = 11

OpenStudy (nubeer):

subsitute 5 in place of x.

OpenStudy (anonymous):

thankyou guys